OpenStudy (anonymous):

Solve the equation. Check both solutions and only write the real solution. sqrt{6m-5}=m A. {1} B. {5} C. {1, 5} D. no real solution

OpenStudy (anonymous):

\[\sqrt{6m-5}=m \implies m^2-6m+5=0=(m-5)(m-1) \implies m=5~or~m=1\]